OpenStudy (anonymous):

Log base b of x = 2 - a + log base b ((a^2 * b^a)/b^2)

OpenStudy (ash2326):

\[\log_b x= 2-a + \log_b \frac{a^2b^a}{b^2}\] Let's bring the log term on the right hand side to left hand side \[\log_b x- \log_b \frac{a^2b^a}{b^2}= 2-a \] We know \[\log_b y- \log_b z= \log _b \frac{y}{z}\] so we get \[\log_b \frac{xb^2}{a^2b^a}= 2-a\] raise both sides to the power of b \[\large b^{\log_b \frac{xb^2}{a^2b^a}}= b^{2-a}\] We know \[\large b^{\log_b y}=y\] so we get here \[\frac{xb^2}{a^2b^a}=b^{2-a}\] Now we have \[\frac{xb^{2-a}}{a^2}=b^{2-a}\] \[\frac{x\cancel{b^{2-a}}}{a^2}=\cancel{b^{2-a}} 1\] we get \[x=a^2\]